OpenStudy (sierra_grey):

The ratio of two numbers is 3 to 5. Their sum is 136. Find the two numbers. how do I write it into an equation?

OpenStudy (tkhunny):

3:5 ==> 3+5 = 8 Hmmm... Multiply by 136/8 = 17 3*17:5*17 ==> 51:85 ==> ?? Did that do it?

OpenStudy (anonymous):

3x+5x=136